A 92.2 g blue jay flies through the trees in order to land at a bird feeder. If it flies with a constant speed of 24.0 km/h and makes a tight horizontal turn with a radius of 3.50 m, how much centripetal force must its wings exert to

make the turn?


1
Expert's answer
2022-01-24T11:25:47-0500

Gives

We know that

Central petal force

M=92.2g

V=25km/h

r=3.50m

"F=\frac{MV^2}{r}"

"F=\frac{92.2\times10^{-3}\times{\frac{24^2\times1000^2}{3600^2}}}{3.50}"

F=1.170N